OpenStudy (anonymous):

What is the area of a sector with radius equal to 8 and measure of arc equal to 300°? 3.3 13.3 53.3

hero (hero):

\[\frac{x}{360} = \frac{y}{\pi r^2}\] x = measure of sector = 300 y = area of sector \[\frac{300}{360} = \frac{y}{\pi (8)^2}\] Solve for y

OpenStudy (anonymous):

360

hero (hero):

@jlinnartz, you have to solve for y

OpenStudy (anonymous):

13

OpenStudy (anonymous):

@Hero

OpenStudy (ivettef365):

I don't see the answer in your choices ?

OpenStudy (anonymous):

The possible answers are 3.3 pi 13.3 pi 53.3 pi

hero (hero):

I'm with @ivettef365

OpenStudy (anonymous):

Well that's what I'm given

OpenStudy (ivettef365):

oh, ok, you have it as pi

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

now I see the answer, ????

OpenStudy (ivettef365):

@Hero gave you the formula, so multiply 64 x 300 then divide by 360

OpenStudy (anonymous):

53.3

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

actually it is 53.33pi

OpenStudy (anonymous):

yes thank you

hero (hero):

So that's what happened...